OpenStudy (anonymous):

Don is drawing a map with a scale of 1 in : 7 mi. He knows that the distance from his house to Fort Collins is 21 miles. How far apart should Don locate the two locations on his map?

OpenStudy (anonymous):

So, you have a map scale of 1 in : 7 mi (which is also 1: 443520) You can treat this like an algebra equation (see, there really *IS* a use for it!) (1) 1 in = 7 mi We have units of 7 mi and 21 mi, and we know that 7 is a factor of 21. 3 is the other factor. Thus: (2) 7 * 3 = 21 Multiply both sides by 3 (3) 3 * 1 in = 3 * 7 mi Simplify (4) 3 in = 21 mi Don should place the locations 3 inches apart on the map to indicate 21 miles in the real world.